Perform the indicated operations

(-1/5)3 + {(-1/3)2 + 1/12(1/6 - (-1/4)3)}

Answer by ptaylor(1894)

Perform the indicated operations
(-1/5)3 + {(-1/3)2 + 1/12(1/6 - (-1/4)3)}
Ok. We will start getting rid of the parens and we start with the innermost parens/brackets first:
Start with -(-1/4)3 and we get
(-1/5)3 + {(-1/3)2 + 1/12(1/6 +3/4)}
Next, we'll add (1/6+3/4) or (2/12+9/12)=11/12 so we now have
(-1/5)3 + {(-1/3)2 + 1/12(11/12)}

Now, we will multiply +1/12(11/12)=11/144 so:
(-1/5)3 + {(-1/3)2 + 11/144}
Next, (-1/3)2=-2/3
(-1/5)3 + {-2/3 + 11/144}
Now, we'll do (-1/5)3=-3/5
-3/5 + {-2/3 + 11/144}
Next, add -2/3+11/144=-96/144+11/144=-85/144
-3/5-85/144
Finally, we will add -3/5-85/144
First, multiply each term by 720/720 to get a common denominator
(-3/5)(720/720)=-432/720
and
(-85/144)(720/720)=-425/720
so
-432/720-425/720=-857/720 or 1.1903--------------------answer

Hope this helps-------ptaylor